# Differentiable Room Acoustic Rendering with Multi-View Vision Priors

### v1

@@ Abstract
An immersive acoustic experience enabled by spatial audio is just as crucial as the visual aspect in creating realistic virtual environments. However, existing methods for room impulse response estimation rely either on data-demanding learning-based models or computationally expensive physics-based modeling. In this work, we introduce Audio-Visual Differentiable Room Acoustic Rendering (AV-DAR), a framework that leverages visual cues extracted from multi-view images and acoustic beam tracing for physics-based room acoustic rendering. Experiments across six real-world environments from two datasets demonstrate that our multimodal, physics-based approach is efficient, interpretable, and accurate, significantly outperforming a series of prior methods. Notably, on the Real Acoustic Field dataset, AV-DAR achieves comparable performance to models trained on 10 times more data while delivering relative gains ranging from 16.6% to 50.9% when trained at the same scale.
@@ 1 Introduction
Existing methods for estimating RIRs generally fall into two broad categories: learning-based and physics-based. Learning-based approaches [37, 39, 48, 58, 5] treat RIR estimation as a regression task trained on densely measured ground-truth RIRs. Although effective for fitting simulated scenes and offering efficient inference, they can be difficult to deploy in real-world settings due to high data requirements and the lack of physically grounded guarantees. In contrast, physics-based approaches [33, 64] rely on explicit acoustic models such as the image-source method [64] or volume rendering [33]. However, both are impractical for large scenes with sparse training samples: the image-source method requires significant computation to enumerate reflection paths and becomes prohibitively expensive in complex environments, while volume rendering demands extensive 3D samples for RIR reconstruction and considerable training data to learn a neural acoustic field.
Our key insight is that while sound fundamentally differs from light—traveling more slowly and exhibiting time-of-arrival variations—both are influenced by the same geometry and surface material properties within a given environment. As shown in Fig. 1, a surface region’s acoustic property often strongly correlates with its visual appearance, as both stem from the same underlying material characteristics. For example, smooth, hard surfaces like glass tend to reflect high-frequency sound, whereas rough, deformable materials such as carpets primarily absorb high-frequency components and reflect lower-frequency waves. Leveraging visual observations to estimate surface acoustic properties in a physics-based differentiable framework could potentially enable more accurate and efficient learning of room acoustic parameters.
@@
Experiments on six real-world environments [13, 64] show that our approach significantly outperforms both learning-based and physics-based baselines. For example, on the Real Acoustic Field dataset [13], our model achieves comparable performance to existing methods trained on roughly 10$\times$ RIR measurements while delivering 16.6% to 50.9% improvement when trained at the same scale. Moreover, qualitative analysis confirms that our method learns material-specific acoustic reflection responses that closely align with the visual characteristics of the scene.
@@ 2 Related Work
Learning-Based RIR Prediction. A growing body of work leverages machine learning techniques to approximate room impulse responses (RIRs) directly. Early methods often rely on large training sets of measured or simulated RIRs to train neural networks capable of predicting new RIRs at new positions. For instance, Ratnarajah et al. [50] use a generative adversarial network (GAN) to synthesize RIRs, while later efforts incorporate scene meshes [49] or visual signals [48] to condition RIR generation. Continuous implicit neural representations have also been explored to model high-fidelity acoustic fields within individual scenes [37, 38, 39, 51, 58, 5]. However, unlike our approach, which can learn from sparse RIR measurments, these methods often rely on dense measurements and/or simulated data, requiring up to 1,000 measured RIRs in order to accurately interpolate to new positions within the same environment [51].
@@ 3.1 Preliminaries
$\widehat{\mathrm{RIR}}(\mathbf{x}_{a},\mathbf{x}_{b},\mathbf{p}_{a},t).\vspace{-0.05in}$ (1)
@@
$\bigl{\{}\{I^{(i)},\pi,P^{(i)}\}{\Big{|}}\;i=1,\cdots,N_{c}\bigr{\}}.\vspace{-0.05in}$ (2)
Our motivation of using multi-view images arises from the fact taht, while RIR measurements encode the final results of acoustic wave propagation—including direct sound, early reflections, and late reverberations—they make it difficult to directly infer the underlying surface reflection properties. Multi-view images, on the other hand, capture vsiual material and geometric cues that are essential for predicting plausible reflection responses, effectively complementing the acoustic modality.
@@
$h_{b}(t)=h(t)\;*\;\widehat{\mathrm{RIR}}(\mathbf{x}_{a},\mathbf{x}_{b},\mathbf{p}_{a},t),\vspace{-0.05in}$ (3)
@@ 3.2 Overview of the AV-DAR Framework
$\widehat{\mathrm{RIR}}(t)=\sum_{\tau}s(\tau;\Theta_{1})\cdot R(t-\tau;\Theta_{2})+r(t;\Theta_{3}),\vspace{-0.1in}$ (4)
@@ 3.3 Acoustic Beam Tracing
$\mathcal{P}(\mathbf{x}_{a},\mathbf{x}_{b})=\{\tilde{\mathbf{x}}_{k}\}_{k=1}^{N}.\vspace{-0.05in}$ (5)
@@
$\prod_{\mathbf{x}_{j}\in\tilde{\mathbf{x}}}\mathrm{Refl}(\mathbf{x}_{j})[f]\cdot\mathbf{D}_{\tilde{x}}[f].\vspace{-0.05in}$ (6)
@@
$R(t)=\sum_{\tilde{\mathbf{x}}_{k}\in\mathcal{P}}\mathcal{S}_{\tilde{t}_{k}}\left\{\kappa(\tilde{\mathbf{x}}_{k},t)\right\}.$
@@ 3.4 Multi-Scale Reflection Response
In a volumetric beam framework, the contact region between a beam and a surface grows as the beam propagates farther from the source, causing the reflected energy to span a larger patch of the surface. This phenomenon naturally introduces a multi-scale effect: near the source, reflections come from a smaller region; farther away, reflections integrate over a larger area. Due to small apex angle, the physical contact region is considered as an ellipse in most cases. However, our beam tracing algorithms only returns a single hit-point ${\bf x}$ for each reflection. To model the overall reflection response from the area, we approximate it by a small Gaussian distribution around the sampled hit point ${\bf x}$. Concretely, we let ${\bf x^{\prime}}\sim\mathcal{N}({\bf x},\Sigma)$, where $\Sigma$ encodes the elliptical patch size and orientation. Intuitively, $\Sigma$ grow with the travel distance $l$ and depends on the reflection angle $\theta$ and half-apex angle $\varphi$. For the detailed calculation of $\Sigma$, please refer to Supp.
@@
${\rm Refl}(\gamma({\bf x},\Sigma);{\Theta}_{2})\in\mathbb{R}^{F},\vspace{-0.05in}$ (9)
@@ 3.5 Multi-View Vision Feature Encoder
We incorporate multi-view images to guide reflection response prediction. Specifically, we learn a vision feature encoder $\mathrm{F}(\mathbf{x},\Sigma;{\Theta}_{2}^{\prime})$ that captures the appearance information around the surface $\mathbf{x}$. The reflection response is then:
$\mathrm{Refl}\Big{(}\gamma({\bf x},\Sigma),\mathrm{F}({{\bf x},\Sigma});{\Theta}_{2}\Big{)}\in\mathbb{R}^{F}\vspace{-0.05in}$ (12)
@@
$\mathbf{v}_{j}^{\mathrm{raw}}={\sum_{i=1}^{N_{c}}}{\rho}\Bigl{(}{\phi(\gamma({\mathbf{z}_{j}}))^{\top}\psi(\mathbf{v}_{j}^{(i)}\oplus P^{(i)})}+m_{j}^{(i)}\Bigr{)}\cdot\alpha(\mathbf{v}_{j}^{(i)})\vspace{-0.05in}$ (14)
where $\mathbf{v}_{j}^{\mathrm{raw}}$ is the intermediate aggregated feature. The functions $\psi$, $\phi$, and $\alpha$ are linear projections to produce keys, quires, and values, respectively. The mask $m_{j}^{(i)}$ is 00 if $\mathbf{z}_{j}$ is visible in $I^{(i)}$ and $\infty$- ∞ if otherwise. $\gamma(\cdot)$ is a Fourier positional encoding, and $\rho$ denotes the softmax function. Finally, a 2-layer feed-forward network (FFN) refines this raw feature:
@@
$\mathrm{F}(\mathbf{x},\Sigma)=\!\sum_{j=1}^{k}\rho\Bigl{(}g\bigl{(}\phi^{\prime}(\gamma(\mathbf{x},\Sigma))-\psi^{\prime}(\mathbf{v}^{*}_{j})+\delta_{j}\bigr{)}\Bigr{)}\bigl{(}\alpha^{\prime}(\mathbf{v}_{j}^{*})+\delta_{j}\bigr{)},\vspace{-0.05in}$ (17)
@@
$\delta_{j}=\vartheta^{\prime}(\mathbf{x}-\mathbf{x}_{j}^{*}).\vspace{-0.05in}$ (18)
@@ 3.6 Position-Dependent Residual Component
Finally, we introduce the residual component $\mathrm{r}(t;{\Theta}_{3})$, which is responsible for modeling high-order reflections, diffuse reflections, diffraction, and late reverberations. We treat each point on the geometry $\mathbf{x}\in\mathcal{M}$ as a secondary sound source and model the residual IR as the integral of these secondary sources from the listener position $\mathbf{x}_{b}$.
@@
$\mathrm{h}(t)=\epsilon(\mathbf{x},\omega,t,\mathbf{x}_{a},\mathbf{p}_{b};{\Theta}_{3}).$ (19)
@@
[…]
$\approx\sum_{k=1}^{N_{r}}S_{\tau_{k}}\{\epsilon\}(\mathbf{x}^{\prime}_{k},-\omega_{k},t;{\Theta}_{3}).\vspace{-0.1in}$ (21)
@@ 4 Experiments
(dB) $\downarrow$ (dB) $\downarrow$ (%) $\downarrow$ (dB) $\downarrow$ (dB) $\downarrow$ (%$\downarrow$ (dB) $\downarrow$ (dB) $\downarrow$ (%) $\downarrow$ (dB) $\downarrow$ (dB) $\downarrow$ (%) $\downarrow$
@@ 4.1 Experiment Settings
Datasets. We evaluate our method on two real-world datasets: the Real Acoustic Field (RAF) [13] dataset and the Hearing Anything Anywhere (HAA) dataset, which are the only available real-world RIR datasets with accompanying visual capture. The RAF dataset contains densely measured monaural RIRs recorded in two office settings (Empty and Furnished) using tens of thousands of source–listener pairs. We use 0.32 s RIR segments resampled at 16 kHz, following prior work [13, 33]. Each room also has a 3D reconstruction and a dense set of images [65].
@@
Loudness $\text{Loudness Error}=\Bigl{|}10\log_{10}\bigl{(}\frac{E_{\mathrm{pred}}}{E_{\mathrm{gt}}}\bigr{)}\Bigr{|},$ (22)
@@
Implementation Details. For multi-view images, we manually select a subset covering the scene, using 13–30 images for most scenes in the two datasets and 65 images for the RAF Furnished room to assess performance saturation (see Supp. for ablation details). In the HAA dataset, where multi-view images are not provided, we randomly sample cameras from the Polycam reconstructions to render images at 512$\times$512 resolution. Note that Polycam is used only for visual rendering, while beam tracing is performed on the original coarse planar geometry. Additional architectural choices and training details are provided in Supp.
@@ 4.2 Quantitative Results
To fully exploit the dense samples in RAF [13] and evaluate performance at various training scales, we split the original training set (80% of all data) into 9 nested subsets ranging from 0.01% to 100% of the data. The smallest subset contains only 3 samples, while the largest includes approximately 30K samples, with each larger subset including all samples from the smaller ones. All models are evaluated on the original test set from [13] to ensure comparability across scales. Table 1 reports results on the 1% dataset and on our model trained with 0.1% of the data. Notably, our method trained on only 0.1% of the data achieves comparable performance to state-of-the-art baselines trained on 10$\times$ more data. In addition, our method consistently outperforms all baselines when trained with the same amount of data, with improvements ranging from 16.6% (Loudness Error in RAF Empty Scene) to 50.9% (T60 Error in RAF Empty Scene). Figure 3 further demonstrates that our model outperforms existing baselines across all training scales, particularly in few-shot settings (below 3% training data).
Figure 4: Signal spatial distribution visualization. Top two rows: Phase and amplitude maps at 0.6 m wavelength. Bottom row: Loudness heatmap. Our model, trained on only 0.1% of the data, accurately captures source directivity and localization, yielding plausible phase and amplitude distributions, while baseline methods fail to reproduce these patterns even with 10$\times$ training data.

### v2

@@ Abstract
An immersive acoustic experience enabled by spatial audio is just as crucial as the visual aspect in creating realistic virtual environments. However, existing methods for room impulse response estimation rely either on data-demanding learning-based models or computationally expensive physics-based modeling. In this work, we introduce Audio-Visual Differentiable Room Acoustic Rendering (AV-DAR), a framework that leverages visual cues extracted from multi-view images and acoustic beam tracing for physics-based room acoustic rendering. Experiments across six real-world environments from two datasets demonstrate that our multimodal, physics-based approach is efficient, interpretable, and accurate, significantly outperforming a series of prior methods. Notably, on the Real Acoustic Field dataset, AV-DAR achieves comparable performance to models trained on 10 times more data while delivering relative gains ranging from 16.6% to 50.9% when trained at the same scale. Project Page: https://humathe.github.io/avdar/.
@@ 1 Introduction
Existing methods for estimating RIRs generally fall into two broad categories: learning-based and physics-based. Learning-based approaches [39, 42, 51, 61, 5, 41] treat RIR estimation as a regression task trained on densely measured ground-truth RIRs. Although effective for fitting simulated scenes and offering efficient inference, they can be difficult to deploy in real-world settings due to high data requirements and the lack of physically grounded guarantees. In contrast, physics-based approaches [35, 67] rely on explicit acoustic models such as the image-source method [67] or volume rendering [35]. However, both are impractical for large scenes with sparse training samples: the image-source method requires significant computation to enumerate reflection paths and becomes prohibitively expensive in complex environments, while volume rendering demands extensive 3D samples for RIR reconstruction and considerable training data to learn a neural acoustic field.
Our key insight is that while sound fundamentally differs from light—traveling more slowly and exhibiting time-of-arrival variations—both are influenced by the same geometry and surface material properties within a given scene. As shown in Fig. 1, a surface region’s acoustic property often strongly correlates with visual appearance due to the same underlying materials. For example, smooth, hard surfaces like glass tend to reflect high-frequency sound, whereas rough, deformable materials such as carpets primarily absorb high-frequency components and reflect lower-frequency waves. Leveraging visual observations to estimate surface acoustic properties in a physics-based differentiable framework could potentially enable more accurate and efficient learning of room acoustic parameters.
@@
[…]
@@ 2 Related Work
Learning-Based RIR Prediction. A growing body of work leverages machine learning techniques to approximate room impulse responses (RIRs) directly. Early methods often rely on large sets of measured or simulated RIRs to train neural networks capable of predicting RIRs at new positions. For instance, Ratnarajah et al[53] use a generative adversarial network (GAN) to synthesize RIRs, while later efforts incorporate scene meshes [52] or visual signals [51, 41] to condition RIR generation. Continuous implicit neural representations have also been explored to model high-fidelity acoustic fields within individual scenes [39, 40, 42, 54, 61, 5]. However, unlike our approach, which can learn from sparse RIR measurments, these methods often rely on dense measurements and/or simulated data, requiring up to 1,000 measured RIRs in order to accurately interpolate to new positions within the same environment [54].
@@ 3.1 Preliminaries
$\widehat{\mathrm{RIR}}(\mathbf{x}_{a},\mathbf{x}_{b},\mathbf{p}_{a},t).\vskip-3.61371pt$ (1)
@@
$\bigl{\{}\{I^{(i)},\pi,P^{(i)}\}{\Big{|}}\;i=1,\cdots,N_{c}\bigr{\}}.\vskip-3.61371pt$ (2)
Our motivation of using multi-view images arises from the fact that, while RIR measurements encode the final results of acoustic wave propagation—including direct sound, early reflections, and late reverberations—they make it difficult to directly infer the underlying surface reflection properties. Multi-view images, on the other hand, capture visual material and geometric cues that are essential for predicting plausible reflection responses, effectively complementing the acoustic modality.
@@
$h_{b}(t)=h(t)\;*\;\widehat{\mathrm{RIR}}(\mathbf{x}_{a},\mathbf{x}_{b},\mathbf{p}_{a},t),\vskip-3.61371pt$ (3)
@@ 3.2 Overview of the AV-DAR Framework
$\widehat{\mathrm{RIR}}(t)=\sum_{\tau}s(\tau;\Theta_{1})\cdot R(t-\tau;\Theta_{2})+r(t;\Theta_{3}),\vskip-7.22743pt$ (4)
@@ 3.3 Acoustic Beam Tracing
$\mathcal{P}(\mathbf{x}_{a},\mathbf{x}_{b})=\{\tilde{\mathbf{x}}_{k}\}_{k=1}^{N}.\vskip-3.61371pt$ (5)
@@
$\prod_{\mathbf{x}_{j}\in\tilde{\mathbf{x}}}\mathrm{Refl}(\mathbf{x}_{j})[f]\cdot\mathbf{D}_{\tilde{x}}[f].\vskip-3.61371pt$ (6)
@@
$R(t)=\sum_{\tilde{\mathbf{x}}_{k}\in\mathcal{P}}\mathcal{S}_{\tilde{t}_{k}}\left\{\kappa(\tilde{\mathbf{x}}_{k},t)\right\}.$ (9)
@@ 3.4 Multi-Scale Reflection Response
In a volumetric beam framework, the contact region between a beam and a surface grows as the beam propagates farther from the source, causing the reflected energy to span a larger patch of the surface. This phenomenon naturally introduces a multi-scale effect: near the source, reflections come from a smaller region; farther away, reflections integrate over a larger area. Due to small apex angle, the physical contact region is considered as an ellipse in most cases. However, our beam tracing algorithms only returns a single hit-point $\bf x$} for each reflection. To model the overall reflection response from the area, we approximate it by a small Gaussian distribution around the sampled hit point $\bf x$}. Concretely, we let ${\bf x^{\prime}}\sim\mathcal{N}({\bf x},\Sigma)$, where $\Sigma$ encodes the elliptical patch size and orientation. Intuitively, $\Sigma$ grows with the travel distance $l$ and depends on the reflection angle $\theta$ and half-apex angle $\varphi$. For the detailed calculation of $\Sigma$, please refer to Supp.
@@
${\rm Refl}(\gamma({\bf x},\Sigma);{\bf\Theta}_{2})\in\mathbb{R}^{F},\vskip-3.61371pt$ (10)
@@ 3.5 Multi-View Vision Feature Encoder
We incorporate multi-view images to guide reflection response prediction. Specifically, we learn a vision feature encoder $\mathrm{F}(\mathbf{x},\Sigma;\mathbf{\Theta}_{2}^{\prime})$ that captures the appearance information around the surface $\mathbf{x}$. The reflection response is then:
$\mathrm{Refl}\Big{(}\gamma({\bf x},\Sigma),\mathrm{F}({{\bf x},\Sigma});\mathbf{\Theta}_{2}\Big{)}\in\mathbb{R}^{F}\vskip-3.61371pt$ (13)
@@
$\mathbf{v}_{j}^{\mathrm{raw}}={\sum_{i=1}^{N_{c}}}{\rho}\Bigl{(}{\phi(\gamma({\mathbf{z}_{j}}))^{\top}\psi(\mathbf{v}_{j}^{(i)}\oplus P^{(i)})}+m_{j}^{(i)}\Bigr{)}\cdot\alpha(\mathbf{v}_{j}^{(i)})\vskip-3.61371pt$ (15)
where $\mathbf{v}_{j}^{\mathrm{raw}}$ is the intermediate aggregated feature. The functions $\psi$, $\phi$, and $\alpha$ are linear projections to produce keys, quires, and values, respectively. The mask $m_{j}^{(i)}$ is 0 if $\mathbf{z}_{j}$ is visible in $I^{(i)}$ and -$\infty$ if otherwise. $\gamma(\cdot)$ is a Fourier positional encoding, and $\rho$ denotes the softmax function. Finally, a 2-layer feed-forward network (FFN) refines this raw feature:
@@
$\mathrm{F}(\mathbf{x},\Sigma)=\!\sum_{j=1}^{k}\rho\Bigl{(}g\bigl{(}\phi^{\prime}(\gamma(\mathbf{x},\Sigma))-\psi^{\prime}(\mathbf{v}^{*}_{j})+\delta_{j}\bigr{)}\Bigr{)}\bigl{(}\alpha^{\prime}(\mathbf{v}_{j}^{*})+\delta_{j}\bigr{)},\vskip-3.61371pt$ (18)
@@
$\delta_{j}=\vartheta^{\prime}(\mathbf{x}-\mathbf{x}_{j}^{*}).\vskip-3.61371pt$ (19)
@@ 3.6 Position-Dependent Residual Component
Finally, we introduce the residual component $\mathrm{r}(t;\mathbf{\Theta}_{3})$, which is responsible for modeling high-order reflections, diffuse reflections, diffraction, and late reverberations. We treat each point on the geometry $\mathbf{x}\in\mathcal{M}$ as a secondary sound source and model the residual IR as the integral of these secondary sources from the listener position $\mathbf{x}_{b}$.
@@
$\mathrm{h}(t)=\epsilon(\mathbf{x},\omega,t,\mathbf{x}_{a},\mathbf{p}_{b};\mathbf{\Theta}_{3}).$ (20)
@@
$\mathrm{r}(t;\mathbf{\Theta}_{3})$ $=\int_{\mathbb{S}^{2}}\mathcal{S}_{\tau}\{\epsilon\}(\mathbf{x}^{\prime}(\omega),-\omega,t;\mathbf{\Theta}_{3})p_{\mathrm{u}}(\omega)d\omega$ (21)
$\approx\sum_{k=1}^{N_{r}}S_{\tau_{k}}\{\epsilon\}(\mathbf{x}^{\prime}_{k},-\omega_{k},t;\mathbf{\Theta}_{3}).\vskip-14.45377pt$ (22)
@@ 4 Experiments
[…]
@@ 4.1 Experiment Settings
Datasets. We evaluate our method on two real-world datasets: the Real Acoustic Field (RAF) [13] dataset and the Hearing Anything Anywhere (HAA) [67] dataset, which are the only available real-world RIR datasets with accompanying visual capture. The RAF dataset contains densely measured monaural RIRs recorded in two office settings (Empty and Furnished) using tens of thousands of source–listener pairs. We use 0.32 s RIR segments resampled at 16 kHz, following prior work [13, 35]. Each room also has a 3D reconstruction and a dense set of images [69].
@@
$\text{Loudness Error}=\Bigl{|}10\log_{10}\bigl{(}\frac{E_{\mathrm{pred}}}{E_{\mathrm{gt}}}\bigr{)}\Bigr{|},$ (23)
@@
Implementation Details. For multi-view images, we manually select a subset covering the scene, using 13–30 images for most scenes in the two datasets and 65 images for the RAF Furnished room to assess performance saturation (see Supp. for ablation details). In the HAA dataset, where multi-view images are not provided, we randomly sample cameras from the Polycam reconstructions to render images at $\times512$ resolution. Note that Polycam is used only for visual rendering, while beam tracing is performed on the original coarse planar geometry. Additional architectural choices and training details are provided in Supp.
@@ 4.2 Quantitative Results
[…]